OpenStudy (anonymous):

The investment portfolio currently contains 3 stocks that have a total value equal to $100,000. The beta is 1.5. The investor is considering investing an additional $50000 in a stock that has a beta of 3. After this is done, what will be the portfolio's new beta

OpenStudy (sweta):

see beta of portfolio is wt avg beta hence, beta of new portfolio = (1.5 * 1 lacs) + (0.5 lacs * 3) whole divided by 1.50 lacs so the ans is 2

OpenStudy (sweta):

Logic to support why wt. avg. beta: Beta is systematic risk in relative terms hence formula will be simply the weighted average, in a diversified portfolio no further diversification possible. Just tried, if wrong please correct.

OpenStudy (anonymous):

Beta of a portfolio is calculated by using this formula \[\beta_{P} = \sum_{i=1}^{n}\beta_{i}weight_{i}\] Now in your case \[\beta_{Security_{1}}=1.5\] and total investment in first security = $ 100000 While second security has a beta of 3 and investment of $ 50000 so what proportion of amout you have invested into first security and second security total investment is $ 150000 First security weight = 0.667 Second Security weight = 0.337 \[\beta_{p} = 0.667 \times1.5 + 0.337 \times 3\] \[\beta_{p} = 1.0005 + 1.011\] \[\beta_{p} = 2.0115\]